# TripleMixer: A 3D Point Cloud Denoising Model for Adverse Weather

### v1

@@ Abstract
LiDAR sensors are crucial for providing high-resolution 3D point cloud data in autonomous driving systems, enabling precise environmental perception. However, real-world adverse weather conditions, such as rain, fog, and snow, introduce significant noise and interference, degrading the reliability of LiDAR data and the performance of downstream tasks like semantic segmentation. Existing datasets often suffer from limited weather diversity and small dataset sizes, which restrict their effectiveness in training robust models. Additionally, current deep learning denoising methods, while effective in certain scenarios, often lack interpretability, complicating the ability to understand and validate their decision-making processes. To overcome these limitations, we introduce two large-scale datasets, Weather-KITTI and Weather-NuScenes, which cover three common adverse weather conditions: rain, fog, and snow. These datasets retain the original LiDAR acquisition information and provide point-level semantic labels for rain, fog, and snow. Furthermore, we propose a novel point cloud denoising model, TripleMixer, comprising three mixer layers: the Geometry Mixer Layer, the Frequency Mixer Layer, and the Channel Mixer Layer. These layers are designed to capture geometric spatial information, extract multi-scale frequency information, and enhance the multi-channel feature information of point clouds, respectively. By leveraging the combined capabilities of these three core mixer layers, the TripleMixer network not only effectively preserves and enhances critical features of point cloud data during the denoising process but also offers strong interpretability. Experiments conducted on the WADS dataset in real-world scenarios, as well as on our proposed Weather-KITTI and Weather-NuScenes datasets, demonstrate that our model achieves state-of-the-art denoising performance. Additionally, our experiments show that integrating the denoising model into existing segmentation frameworks enhances the performance of downstream tasks. These results highlight our model’s potential to significantly improve the accuracy and reliability of autonomous driving systems in complex environments. The datasets and code will be made publicly available at https://github.com/Grandzxw/TripleMixer.
@@
LiDAR Dataset, Weather-KITTI, Weather-Nuscenes, Point Cloud Denosing, Adverse Weather
@@ I Introduction
Figure 1: Visualization of semantic 3D points from the same LiDAR frame across four weather conditions in our proposed Weather-KITTI dataset. Weather noises are marked in red.
LiDAR sensors play a crucial role in autonomous driving systems by emitting laser beams and capturing the reflected signals to generate high-resolution 3D point cloud data, thus providing precise environmental perception[1, 2, 3, 4]. However, the LiDAR data included in current large-scale autonomous driving datasets are typically ”clean”. In real-world scenarios, varying weather conditions, such as rain, fog, and snow, can cause laser beams to scatter, be absorbed, or reflect abnormally, leading to significant noise and interference[5, 6, 7]. This increase in noise not only reduces the reliability of the LiDAR data but also potentially degrades the accuracy of downstream tasks, such as semantic segmentation and object detection, which in turn can affect the overall decision-making capability of autonomous driving systems[8, 9, 10, 11]. Therefore, effectively denoising LiDAR data while preserving its accuracy under diverse weather conditions has become a critical challenge in autonomous driving systems.
To address this challenge, some researchers have specifically developed LIDAR datasets under adverse weather conditions. These datasets can be broadly categorized into two types: real-world datasets, collected under adverse weather conditions, and synthetic datasets, generated through computer simulations[12, 13, 14]. Real-world datasets often face challenges related to the complexity of weather annotation, high data collection costs, and small dataset sizes, which restrict their applicability across a wide range of weather conditions. On the other hand, synthetic datasets allow researchers to create extreme weather scenarios under controlled conditions, enabling more comprehensive testing of system reliability. However, existing synthetic datasets suffer from limitations such as limited weather types, the loss of original LiDAR acquisition information and a lack of precise point-level weather labels.
In addition to these datasets, several denoising methods for point clouds have been proposed, which can be broadly categorized into statistical approaches and deep learning methods[12, 13]. Statistical approaches filter noise by analyzing the distribution patterns of point clouds. However, despite their effectiveness on smaller datasets, they often require substantial computational resources, especially when dealing with large-scale point clouds[15, 16, 17, 18, 19]; In contrast, deep learning methods leverage the powerful feature-learning capabilities of neural networks to directly learn noise patterns from input point clouds. Existing deep learning denoising networks are typically further classified into two categories based on the input format of the point clouds: range image-based methods and 3D point cloud-based methods[20, 21, 22, 23]. The former involves projecting 3D point cloud data onto a 2D range images, but this projection process may lead to a loss of spatial information, potentially affecting the final denoising performance, particularly when dealing with complex point cloud data. The latter directly processes raw 3D point cloud data, better preserving the spatial structure and geometric information, which results in more effective denoising. However, while these deep learning methods have achieved promising results in denoising tasks, they often lack interpretability.
To address the challenges posed by existing datasets and denoising methods, this paper presents two key solutions. First, we introduce two large-scale adverse weather datasets, Weather-KITTI and Weather-NuScenes, which cover three common adverse weather conditions, including rain, fog, and snow. These datasets retain the original LiDAR acquisition information, such as intensity, and provide point-level labels for rain, fog, and snow. Our datasets not only better reflect the performance of LiDAR sensors under different weather conditions in the real world but also provide researchers with more precise and detailed annotations, supporting deeper algorithm research and model optimization. This design aims to provide robust and reliable data support for improving the robustness and reliability of autonomous driving systems under complex weather conditions. Second, we propose a novel point cloud denoising model, the TripleMixer network, which consists of three mixer layers: the Geometry Mixer Layer, Frequency Mixer Layer, and Channel Mixer Layer. The Geometry Mixer Layer focuses on capturing the geometric spatial information of point clouds by analyzing and mixing the geometric relationships of neighboring points, thereby preserving local structural features and important geometric details during the denoising process. The Frequency Mixer Layer is dedicated to extracting frequency information from point clouds by employing multi-scale analysis. This process effectively learns the distinct spectral features from low-frequency components, which represent the global structure, and from high-frequency components, which contain both details and noise, improving the accuracy of point cloud interpretation in complex environments. The Channel Mixer Layer enhances the multi-channel feature information of point clouds by facilitating information exchange and integration between different feature channels, capturing contextual relationships between features, and further improving denoising accuracy. The collaborative effect of these three core mixer layers enables the TripleMixer network to not only preserve and enhance critical features of point cloud data during the denoising process but also provides strong interpretability. Experiments conducted on the WADS dataset[17] in real-world scenarios, as well as on our proposed weather-KITTI and Weather-Nuscene datasets, demonstrate that our model achieves state-of-the-art denoising performance. Additionally, when applying our denoising model to downstream semantic segmentation tasks, experimental results indicate that integrating our denoising model into existing segmentation models can notably improve their segmentation performance. Our main contributions are summarized as follows:
  • We introduce two large-scale datasets, Weather-KITTI and Weather-NuScenes, specifically designed for autonomous driving in three adverse weather conditions: rain, fog, and snow. These datasets retain the original LiDAR acquisition information, including intensity, and provide precise point-level weather labels. We believe that these datasets will support efforts to improve the robustness and reliability of autonomous driving systems under complex weather conditions.
  • We propose a novel point cloud denoising model, the TripleMixer network, which consists of three distinct mixer layers: the Geometry Mixer Layer, the Frequency Mixer Layer, and the Channel Mixer Layer. These layers effectively preserve geometric spatial information, separate frequency components, and enhance multi-channel feature integration. Our TripleMixer network not only preserves and enhances critical features of point cloud data during the denoising process but also provides strong interpretability.
  • Extensive experiments conducted on both existing and newly proposed datasets demonstrate the state-of-the-art denoising performance of the proposed TripleMixer model, and show that integrating our denoising model into existing semantic segmentation frameworks leads to notable improvements in downstream tasks. These findings highlight the practical applicability and effectiveness of the proposed model in enhancing the robustness and reliability of autonomous driving systems in real-world scenarios.
@@ II Related Work
II-A Adverse Weather Datasets
Weather changes are common in the physical world of autonomous driving. Recent datasets investigating LiDAR performance in adverse weather conditions are generally divided into two categories: real-world recordings and synthetic point clouds. Existing real-world datasets face challenges with complex manual weather labeling and limited scale. Meanwhile, current synthetic datasets often fail to retain the original characteristics of LiDAR.
Real-world datasets are typically collected either outdoors or in weather chambers. The DENSE dataset[24] was recorded within fog chambers, which allow for full control over weather conditions and surroundings. However, despite this level of control, these chambers do not adequately reflect real-world conditions. The CADC dataset[25] was recorded during the wintertime in adverse Canadian weather with the VLP32 LiDAR sensors. The dataset has 3D bounding box annotations available, but does not come with weather-related labels and it only covers snowfall weather conditions. The Boreas dataset[26] was collected by driving a repeated route under adverse weather conditions such as rain and falling snow, using a 128-channel LiDAR. It provides 3D bounding box annotations for 3D object detection, but also does not offer weather-related labels. The WADS dataset[17] is an extreme weather dataset from open-world scenarios and includes point-wise weather labels. However, due to the complexity of manually labeling snow noise points individually, it features a limited scale and few scenarios.
Synthetic datasets are collected by augmenting existing datasets. The SnowKITTI dataset[21] is a synthetic dataset based on the KITTI odometry benchmark dataset[27]. However, it only contains data from snowfall weather conditions and has altered the original KITTI point cloud format, with each point cloud frame formatted as 1048×64. Additionally, it has lost the intensity information of each point cloud. The SMART-Degradation dataset[28] provides additional raw data collected by a 32-channel 3D LiDAR and synthesized rainy LiDAR scans. However, it does not provide point-wise weather labels for the rainy LiDAR scans, making it challenging to apply to point cloud perception tasks.
II-B Point Cloud Denoising Method
Adverse weather conditions can impair the effectiveness of LiDAR sensors by introducing noises into their measurements. These noises can significantly degrade the performance of tasks such as perception and mapping[29]. Consequently, denoising point clouds under such conditions is essential to restore measurement clarity. The emerging field of denoising LiDAR point clouds in adverse weather conditions has been explored in only a few studies so far. The primary methodologies employed include statistical approaches and deep learning approaches.
Statistical approaches filter noise by analyzing the distribution patterns of point clouds. Statistical Outlier Removal (SOR) and Radius outlier removal (ROR) [16] filter out noise by analyzing the density of the nearest neighbor points. SOR removes a noisy point by comparing the mean and standard deviation of the distance to its k-nearest neighbors. ROR processes each point in the point cloud, counting the number of neighbors within a specified search radius to remove noise points. Dynamic ROR (DROR) [30] enhances the ROR filter by dynamically adjusting the search radius. Dynamic SOR (DSOR) [17] is an extension of SOR filter designed to address the non-uniform density within point clouds. It dynamically calculates a local distance threshold to remove noise points. Low-Intensity Outlier Removal (LIOR)[19] and Dynamic Distance–Intensity Outlier Removal (DDIOR)[18] integrates the distance and intensity of points based on the systematic and accurate analysis of data characteristics in snowy weather. Conditional Random Fields Outlier Removal (CRFOR)[31] proposes a de-snowing formulation with Conditional Random Fields and Spatio-Temporal information. Since statistical approaches typically remove noise through iterative algorithms, they are computationally expensive.
Contrary to statistical approaches, deep learning approaches became popular due to the model’s abilities to directly understand the characteristics of weather-induced noise in point cloud. Some general-purpose point cloud semantic segmentation approaches [32, 33, 34] have been adapted to denoise point clouds affected by adverse weather. Although they have achieved good performance, they are not specifically designed for denoising. Existing denoising networks are typically categorized based on the input format of the point cloud: range images and raw 3D point clouds. WeatherNet[20] uses modified LiLaBlock[35] to filter out adverse weather effects in point cloud data. 4DenoiseNet[21] proposes a k-nearest neighbor convolution kernel that captures spatial-temporal information, and with a attention mechanism. However, its limitation lies in the requirement for LiDAR scans with temporal continuity. LiSnowNet[36] combines the Fast Fourier Transform (FFT) with Convolutional Neural Networks (CNNs) for unsupervised de-noising to remove snow from LiDAR point clouds. The denoising networks mentioned above all use range projection images as inputs. While range projection methods usually benefit from fast computations, they face limitations such as projection errors and loss of spatial information when converting 3D point clouds into 2D range images. 3D-OutDet[22] takes raw 3D point clouds as inputs and develops a NeighborHood (NH) Convolution for unorganized scattered 3D point clouds, specifically targeting the removal of snowflakes.
@@ II-B Point Cloud Denoising Method
Table I: Comparison of the Proposed Adverse Weather Dataset with Existing Public Datasets. Including: Year, LiDAR Parameters, Number of Dataset Frames, Presence of Semantic Labels, Types of adverse Weather Covered,
Inclusion of LiDAR Intensity Features, and Presence of Point-Wise Weather Labels.
@@
CADC[25] 2021 32 channel 7K frames bounding boxes
DENSE[24] 2020 32 & 64 channel 13K frames bounding boxes
Boreas[26] 2023 128 channel 7K frames bounding boxes
SMART-Degradation[28] 2021 32 channel 1K frames -
Oxford RobotCar[37] 2020 32 channel 11K frames -
SemanticKITTI[38] 2019 64 channel 2K frames point-wise
nuScenes-lidarseg[39] 2020 32 channel 28K frames point-wise
SnowKITTI[21] 2023 64 channel 43K frames -
@@
III Datasets Generation
@@
As introduced in the previous section, current adverse weather datasets face several challenges: Firstly, there is an absence of point-wise semantic labels and weather labels; Secondly, manually annotating weather conditions in large-scale, real-world datasets is highly complex; Thirdly, existing synthetic datasets lack the original LiDAR acquisition information such as intensity, which impacts the performance of LiDAR downstream perception tasks. Building on this, we propose our synthetic adverse weather datasets, named Weather-KITTI and Weather-NuScenes, which are based on the SemanticKITTI[38] and nuScenes-lidarseg[39] datasets, respectively. To simulate three adverse weather conditions: snow, fog, and rain, we employ the LiDAR Snowfall Simulation (LSS) [40] for snow, the LiDAR Fog Simulation (LFS) [41] for fog, and the LiDAR Light Scattering Augmentation (LISA) [42] for rainy conditions. Table I presents the characteristics of the proposed dataset and existing publicly datasets. Compared to these existing adverse weather datasets, our proposed dataset provides a larger scale, covers a broader range of weather scenarios, offers complete original point cloud information, and has the capability to provide point-wise semantic and weather labels.
III-A Adverse Weather Simulator
LiDAR sensors emit laser pulses and measure travel distance based on the echoes received from targets, thereby capturing the precise shape of the targets[43]. However, the LiDAR pulse system is susceptible to interference from scattering media such as particles of rain, fog, and snow, which can lead to shifts in measured distances, variations in echo intensity, and missing points, among other issues[44, 45, 46].
The received signal power at a LiDAR’s receiver can typically be described using a linear system model[43], which is valid for non-elastic scattering. In this model, the range-dependent received signal power $P_{R}$ is modeled as the time-wise convolution between the time-dependent transmitted signal power $P_{T}$ and the time-dependent impulse response $H$ of the environment:
[…]
where $C_{A}$ is a system constant independent of time $t$, $R_{0}$ is the object distance and $c$ is the speed of light.
1) Snow Model: In this work, we follow [40] to simulate these snow-corrupted data. This physically-based method samples snow particles in the 2D space and modify the measurement for each LiDAR beam in accordance with the induced geometry. For each point $p(x,y,z,i)$ from LiDARs, we calculate its snow-attenuated point $p_{\text{snow}}$ as follows:
$R^{*}=\arg\max\left({P}_{R}\right)-c\frac{\tau_{H}}{2},\$ (2)
$i_{\text{snow}}=P_{R}+i_{\max}f_{s}\left|f_{o}-\left(1-\frac{R^{*}}{R_{\max}}\right)\right|^{2},\$ (3)
${{p_{\text{snow}}}}=\left(\frac{R^{*}}{R_{0}}\cdot(x,y,z),i_{\text{snow}}\right),$ (4)
where $\tau_{H}$ is the LiDAR sensor’s half-power pulse width. $R^{*}$ and ${R_{\max}}$ represent the distances to the object at maximum received power $P_{R}$ and the furthest detectable object, respectively. $f_{s}$ is the focal slope and $f_{o}$ is the focal offset of the LiDAR sensor. As the snowfall rate $r_{s}$ increases, so does the number of particles in a point cloud frame, leading to denser snow particle accumulations in heavier snowfalls. Similar to [17], we define point clouds with changed spatial positions as the snow class.
2) Fog Model: The physically-based fog simulation[41] is employed to transform clear LiDAR point clouds into their foggy counterparts. For each original point cloud, we calculate its fog-attenuated point $p_{\text{fog}}$ as follows:
$i_{\text{hard }}={i}\cdot\exp({-2\alpha R_{0}}),$ (5)
$i_{\text{soft}}={i}\cdot\frac{R_{0}^{2}}{\beta_{0}}\beta\cdot{i}_{tmp},$ (6)
${{p_{\text{fog}}}}=\left\{\begin{array}[]{ll}\left(s\cdot n\cdot(x,y,z),{i_{\text{soft }}}\right),&\text{ if }{i_{\text{soft }}}>{i_{\text{hard }}}\\ \left(x,y,z,{i_{\text{hard }}}\right),&\text{ else. }\end{array}\right.$ (7)
where $\alpha$ is the attenuation coefficient, $\beta$ denotes the fog backscattering coefficient, $\beta_{0}$ describes the differential reflectivity of the target. $i_{\text{soft }}$ and $i_{\text{hard }}$ represent the received responses for the soft target term and the hard target term, respectively. $s$ denotes scaling factor and $n$ is noise factor. ${i}_{tmp}$ is the maximum response of the soft-target term. As $\beta$ increase, the fog rate rises, and the number of fog particles in the point cloud correspondingly increases. We similarly map point clouds with changes in three-dimensional spatial positions as the fog class.
2) Rain Model: We employ the physics-based simulation model[42] to augment LiDAR scans for rain condition, which is based on a hybrid Monte-Carlo based approach. Specifically, the rain-distorted point cloud ${{p_{\text{rain}}}}$ is calculated as follows:
[…]
${i_{\text{rain}}}={i}\cdot\exp^{-2\alpha R_{\text{max }}}\cdot\min\left\{\left(\frac{D_{\text{max }}}{D_{b}\left(R_{\text{max }}\right)}\right)^{2},1\right\},$ (9)
${{p_{\text{rain}}}}=\left(\gamma\cdot(x,y,z),i_{\text{rain}}\right),$ (10)
where $\alpha$ is the extinction coefficient, $D$ is the particle diameter and $q_{e}$ is the extinction efficiency. $N$ denotes the particle density function. $D_{b}$ represents the beam diameter at the object distance $R_{\text{max }}$ and $\gamma$ is the rain-attenuated coefficient. Similarly, as the rain rate $r_{r}$ increases, the number of rain particles in a point cloud frame also correspondingly increases. We map point clouds with a rain-attenuated coefficient $\gamma$ not equal to 1 as the rain class.
@@ III-A Adverse Weather Simulator
Figure 3: Semantic distribution of 3D points for all sequences in Weather-KITTI and Weather-NuScenes.
Table II: Details of Weather Simulations. Each Weather
Simulation Is Categorized Into Light, Moderate,
and Heavy Levels Based on Different
Weather Parameters.
Weather Intensity Description

### v2

@@ Abstract
Adverse weather conditions such as snow, fog, and rain pose significant challenges to LiDAR-based perception models by introducing noise and corrupting point cloud measurements. To address this issue, we propose TripleMixer, a robust and efficient point cloud denoising network that integrates spatial, frequency, and channel-wise processing through three specialized mixer modules. TripleMixer effectively suppresses high-frequency noise while preserving essential geometric structures and can be seamlessly deployed as a plug-and-play module within existing LiDAR perception pipelines. To support the development and evaluation of denoising methods, we construct two large-scale simulated datasets, Weather-KITTI and Weather-NuScenes, covering diverse weather scenarios with dense point-wise semantic and noise annotations. Based on these datasets, we establish four benchmarks: Denoising, Semantic Segmentation (SS), Place Recognition (PR), and Object Detection (OD). These benchmarks enable systematic evaluation of denoising generalization, transferability, and downstream impact under both simulated and real-world adverse weather conditions. Extensive experiments demonstrate that TripleMixer achieves state-of-the-art denoising performance and yields substantial improvements across all downstream tasks without requiring retraining. Our results highlight the potential of denoising as a task-agnostic preprocessing strategy to enhance LiDAR robustness in real-world autonomous driving applications.
@@
Adverse Weather Dataset, Point Cloud Denosing, Robust LiDAR perception, Plug-and-Play Network
@@ I Introduction
Figure 1: Visualization of semantic 3D points from the same LiDAR frame across four weather conditions in our proposed Weather-KITTI dataset. Weather noises are marked in red and loss points are marked in pink.
Robust LiDAR perception [1, 2] in adverse weather remains a critical challenge for autonomous driving systems. Weather-induced degradations such as snow, fog, and rain can cause LiDAR beams to scatter, reflect, or be absorbed abnormally, leading to corrupted point cloud measurements with substantial noise and missing points [3]. These degradations compromise the reliability of 3D sensing and degrade the performance of downstream perception tasks [4, 5, 6, 7, 8], including semantic segmentation, place recognition, and object detection.
To address this issue, recent research has explored two broad categories of solutions: domain transfer methods [9, 10, 11] and data preprocessing via denoising [12, 13, 14]. Domain transfer methods, including domain adaptation and domain generalization techniques, aim to improve model robustness by adapting features across clean and adverse-weather domains. However, they typically require access to target domain data, involve task-specific retraining, and often lack generalizability across different perception tasks. In contrast, denoising serves as a task-agnostic preprocessing step that can be seamlessly integrated into existing LiDAR perception pipelines without the need for retraining.
Existing denoising approaches can be broadly categorized into statistical and learning-based methods [15, 16]. Statistical methods [17, 18, 19] rely on local geometric heuristics, but their iterative filtering procedures make them computationally inefficient and poorly scalable to large-scale point clouds. Learning-based methods typically follow one of two paradigms: projecting 3D point clouds onto 2D range images [20, 21] or directly processing raw point clouds [14]. Most existing models exhibit limited ability to capture multi‑scale spatial structural information and lack frequency‑aware modeling to distinguish meaningful structures from high‑frequency noise, resulting in suboptimal denoising performance under adverse weather conditions.
To overcome these challenges, we propose TripleMixer, a robust and efficient point cloud denoising architecture tailored for adverse weather conditions. TripleMixer consists of three complementary components: (1) a Geometry Mixer (GMX) that captures local geometric structures from raw 3D points while preserving fine-grained spatial details; (2) a Frequency Mixer (FMX) that performs multi-scale frequency decomposition across three orthogonal 2D projections to effectively separate structural information from high-frequency noise; and (3) a Channel Mixer (CMX) that facilitates inter-channel feature refinement to enhance contextual representation. This design enables TripleMixer to suppress noise while preserving critical geometric features, striking an effective balance between performance, efficiency, and interpretability.
To advance the research community’s capabilities in robust LiDAR perception under adverse weather, and to support the development and rigorous evaluation of TripleMixer, we construct two large-scale simulated datasets: Weather-KITTI and Weather-NuScenes. As shown in Table I, our datasets contain over 200K LiDAR frames, substantially exceeding the scale of existing benchmarks, which typically include fewer than 50K frames. Moreover, our datasets preserve the characteristics of real LiDAR acquisitions while covering multiple adverse weather types (snow, fog, and rain). They also provide both point-wise semantic labels and weather annotations, which are lacking in most existing datasets. Built upon these datasets, we establish four benchmarks: (1) Denoising benchmark for evaluating denoising performance in isolation, and three downstream perception benchmarks under real-world adverse weather conditions: (2) Semantic Segmentation (SS), (3) Place Recognition (PR), and (4) Object Detection (OD). Our main contributions are summarized as follows:
  • We propose TripleMixer, a plug-and-play point cloud denoising network that integrates spatial, frequency, and channel-wise processing through three specialized mixer layers. TripleMixer enables interpretable and robust denoising under adverse weather conditions, and can be seamlessly integrated into existing LiDAR perception pipelines to enhance their robustness.
  • We introduce Weather KITTI and Weather NuScenes, two large-scale simulated LiDAR datasets encompassing diverse adverse weather conditions. Based on these datasets, we establish four benchmarks: Denoising, SS, PR, and OD. These benchmarks support comprehensive evaluation and provide a solid foundation for developing robust denoising and 3D perception models under both simulated and real-world adverse weather conditions.
  • Extensive experiments on both all benchmarks demonstrate that TripleMixer achieves state-of-the-art denoising performance. Furthermore, when used as a preprocessing module without retraining downstream models, TripleMixer significantly improves performance on SS, PR, and OD tasks under real-world adverse weather, achieving relative accuracy gains of up to 20%, 56%, and 16%, respectively.
@@ II Related Work
II-A Robust LiDAR Perception
Adverse weather conditions pose significant challenges to LiDAR-based perception by introducing severe noise and degradation into point cloud measurements [22]. Recent efforts to address this issue primarily fall into two categories: domain transfer and data preprocessing through denoising.
Domain transfer methods include domain adaptation [9, 23, 11], which leverages labeled or unlabeled target-domain (i.e., adverse weather) data to align feature distributions via retraining or fine-tuning, and domain generalization [24, 10], which enhances model robustness through data augmentation or domain-invariant representation learning. However, these approaches typically require access to target-domain data, involve task-specific retraining, and often struggle to generalize across diverse perception tasks.
In contrast, denoising offers a task-agnostic and modular preprocessing solution that can be seamlessly integrated into existing LiDAR perception pipelines. Existing denoising approaches can be broadly categorized into statistical and deep learning-based methods. Statistical methods, such as SOR and ROR [17], rely on local geometric heuristics, while dynamic variants like DROR [25] and DSOR [18], as well as intensity-aware filters such as LIOR [19] and DDIOR [26], offer improved resilience to weather-induced noise. However, their iterative nature makes them computationally expensive and limits scalability. Deep learning-based approaches learn complex noise patterns from data and are typically grouped by input modality: 2D range images or raw 3D point clouds. Range image-based methods, such as WeatherNet [20], LiSnowNet [27], and 4DenoiseNet [21], leverage CNNs or FFTs to process projected range representations, but inherently compromise the fidelity of 3D spatial structure. In contrast, point cloud-based models like 3D-OutDet [14] directly operate on raw point clouds using neighborhood-based convolutions, preserving geometric detail but introducing significant runtime overhead due to $k$-NN preprocessing. Despite recent progress, existing methods remain limited in their capacity to capture multi-scale spatial structure and generally lack frequency-aware modeling, which is critical for distinguishing meaningful features from high-frequency noise.
II-B Adverse Weather Datasets and Benchmarks
Existing adverse-weather datasets for LiDAR perception can be broadly categorized into two types: real-world datasets and simulated datasets. Real-world datasets are collected using actual sensors under varying environmental conditions, providing realistic examples of sensor behavior in adverse weather. For instance, the STF dataset [28] offers multimodal recordings with different levels of visibility, while the CADC dataset [29] captures winter driving scenes using a VLP-32 LiDAR and provides 3D bounding box annotations. The WADS dataset [18] focuses on extreme outdoor environments and includes point-wise annotations of weather-induced noise, making it suitable for fine-grained analysis.
Simulated datasets are typically created by applying weather simulators to existing clean-weather LiDAR datasets. These models simulate the physical interactions between LiDAR beams and environmental factors such as rain, fog, and snow. For example, a hybrid Monte Carlo model is employed in [30] to simulate raindrop interference, while fog effects are modeled using data-driven and physically based methods in [31, 32]. Snow simulation is handled in [33] by superimposing reflections from virtual particles. Building on these techniques, datasets such as SnowKITTI [21] and Robo3D [34] augment existing datasets like KITTI [35], SemanticKITTI [36], and nuScenes [37] with weather effects. Some recent benchmarks [38, 39] also introduce synthetic perturbations across multiple sensor modalities to evaluate model robustness.
Overall, real-world datasets offer valuable realism by capturing authentic environmental variations and sensor responses. However, their scale is often constrained by the high cost of manual annotation, and they typically provide limited coverage across diverse weather conditions. In contrast, simulated datasets enable controllable generation of adverse weather effects and can effectively augment existing clean-weather datasets. Nonetheless, most existing simulated datasets are restricted in scale, lack fine-grained weather annotations, and have yet to be systematically validated for their effectiveness in enhancing downstream perception tasks under real‑world adverse weather conditions. To address these limitations, we introduce two large-scale simulated datasets and establish four benchmarks spanning both denoising and downstream LiDAR perception tasks under real-world adverse weather scenarios.
@@ II-B Adverse Weather Datasets and Benchmarks
Table I: Comparison of the Proposed Adverse Weather Dataset with Existing Public Datasets, including Year, LiDAR Parameters, Frame Count, Semantic Labels, Weather Types, Intensity Features, and Point-Wise Weather Noise Labels.
@@
CADC[29] 2021 32 channel 7K frames object-wise
STF[28] 2020 32 & 64 channel 13K frames object-wise
LIBRE[40] 2020 32 & 64 & 128 channel 6K frames object-wise
Boreas[41] 2023 128 channel 7K frames object-wise
SemanticSTF[23] 2023 64 channel 2K frames point-wise
Robo3d-KITTI[34] 2023 64 channel 8K frames object-wise
Robo3d-SemanticKITTI[34] 2023 64 channel 8K frames point-wise
Robo3d-nuScenes[34] 2023 32 channel 12K frames point-wise
Robo3d-WOD[34] 2023 64 channel - - -
SnowKITTI[21] 2023 64 channel 43K frames
@@
  • : SemanticSTF provides weather‑related labels for indiscernible points under adverse weather but lacks separate noise annotations,
    ✓: included, ✗: not included, –: unknown,
III Weather‑KITTI and Weather‑NuScenes Datasets
@@
As introduced earlier, we propose our simulated adverse weather datasets, named Weather-KITTI and Weather-NuScenes, which are based on the SemanticKITTI[36] and nuScenes-lidarseg[37] datasets, respectively. To simulate three adverse weather conditions: snow, fog, and rain, we employ the LiDAR Snowfall Simulation (LSS) [33] for snow, the LiDAR Fog Simulation (LFS) [32] for fog, and the LiDAR Light Scattering Augmentation (LISA) [30] for rainy conditions. These simulation methods produce realistic weather effects, and they have demonstrated that perception models trained on their simulated data exhibit improved robustness under adverse weather conditions. Table I provides a detailed comparison between the proposed dataset and existing publicly available adverse‑weather datasets. Compared to these datasets, our dataset offers a larger scale, covers a broader range of weather scenarios, preserves complete point cloud information, and provides dense point‑wise semantic and weather annotations.
@@
((c))
((d))
Figure 3: Top: Semantic distribution of 3D points for all sequences in Weather-KITTI and Weather-NuScenes. Bottom: Distribution of weather-affected points across semantic classes under snow, fog, and rain. “Retained” denotes the percentage of points in that class unaffected by weather artifacts.
Table II: Details of Weather Simulations.
Weather Severity Description